# Identifying the magnetic genes in fully- and partially-ordered V2Al ( Cr, Mn, Fe, Co, Ni) Heusler alloys

### v1

@@ III.3 Magnetic transition temperature of - and -V2Al
The temperature dependence of magnetic phase transitions in -type VAl compounds differs markedly from that in -type structures. As shown in Fig.3(b), the Curie temperatures in -type compounds vary within a narrower range, from approximately 700 K to 300 K, compared to the broader variation in -type materials. This narrower range reflects the more complex network of magnetic exchange couplings in the inverse-Heusler structure. Specifically, V2CrAl exhibits the highest at about 700 K, while the Curie temperature gradually decreases as changes from Fe to Ni. The magnetic transition temperature of V2MnAl shows a slight anomaly, lying between the values for Co and Ni. Despite the apparent complexity of these trends, they can be rationalized using the proposed microscopic mechanism based on the triangular exchange pathways. As discussed above, the NN exchange coupling involving the -site dominates the magnetic transition behavior. For instance, the relatively large contributes to the higher in V2CrAl. Although is also sizable, the of V2MnAl is significantly lower due to the additional influence of the coupling. In V2CrAl, the ferromagnetic – interactions enhance , whereas in V2MnAl, the coupling is antiferromagnetic, leading to a notable reduction of . In the series Fe, Co, Ni, the gradual decrease of is primarily driven by the weakening of the NN coupling, which diminishes from to 0 meV.
@@
III.4 Effect of chemical disorder on magnetism of partially-ordered V2Al
@@
Our calculations on partially chemically disordered VAl indicate that the NN magnetic exchange coupling remains the dominant factor governing the magnetic transition temperature. Notably, during the transition from the to the -type structure, the 4 site is consistently occupied by V2 atoms, ensuring that the V-X-V magnetic triangular pathway is preserved. Consequently, the microscopic magnetic mechanism in chemically disordered VAl is still primarily governed by the V-X-V triangular pathway and the associated NN magnetic exchange couplings .
III.5 Electronic structures of fully-ordered V2Al
Based on the calculated magnetic ground states of -type VAl in Sec. III.1, the observed complex magnetic orders and non-monotonic variation of total magnetic moments with atomic composition are proposed related to their electronic structures. In -type VAl, the relatively small total magnetic moments of V2CrAl and V2MnAl originate from the anti-parallel alignment of spin moments between the V and sublattices. This behavior is clearly reflected in the calculated density of states (DOS). As shown in Fig. 6(a), the total DOS of -type V2Al exhibits a high value at the Fermi level, indicating itinerant magnetism in these compounds [46, 47, 48]. Notably, several compounds display nearly half-metallic character. As shown in Fig. 6(b), Cr and Mn atoms exhibit pronounced occupied states with sharp peaks at approximately -1 eV and -2 eV in the spin-down channel, while their spin-up channel states are minimally occupied. In contrast, for Fe, Co, and Ni, the PDOS of atoms shows significant occupation in the spin-up channel and comparatively low values in the spin-down channel. These electronic features are consistent with the observed transition in magnetic order from ferrimagnetism ( Cr, Mn) to ferromagnetism ( Fe, Co, Ni).
@@ IV CONCLUSION
In this work, we employed density functional theory calculations combined with Monte Carlo simulations based on Heisenberg magnetic exchange couplings to investigate V2Al alloys and introduced the concept of “magnetic genes”. These magnetic genes govern both the magnetic ground states and the magnetic transition temperatures in fully- and partially-ordered compounds. The underlying physical mechanism arises from the modulation of NN magnetic exchange couplings within the V--V triangular pathways, which is influenced by atomic composition, crystal structure, and the degree of chemical disorder. Using the framework of magnetic genes, the magnetic orders can be rationalized in terms of the V--V triangular pathways formed by NN couplings. Furthermore, the magnetic transition temperatures are predominantly determined by both NN and additional magnetic exchange interactions, irrespective of whether the alloy is fully or partially ordered. We anticipate that the identification of magnetic genes in other Heusler alloys may provide a systematic and predictive approach to understanding and tuning their diverse magnetic properties.
@@ Acknowledgements.
This work is supported by the National Natural Science Foundation of China (Grants No. 52371174 and No. 12204033), the Fundamental Research Funds for the Central Universities (Grant No. FRF-TP-22-097A1), the State Key Lab of Advanced Metals and Materials (Grant No. 2022Z-13), the Young Elite Scientist Sponsorship Program by BAST (Grant No. BYESS2023301) and the High-performance Computing Platform of University of Science and Technology Beijing.References
- Gupta et al. [2023] S. Gupta, S. Chakraborty, V. Bhasin, S. Pakhira, S. Dan, C. Barreteau, J.-C. Crivello, S. N. Jha, M. Avdeev, J.-M. Greneche, D. Bhattacharyya, E. Alleno, and C. Mazumdar, High spin polarization in the disordered quaternary heusler alloy femnvga, Phys. Rev. B 108, 045137 (2023).
- Venkateswara et al. [2023] Y. Venkateswara, J. Nag, S. S. Samatham, A. K. Patel, P. D. Babu, M. R. Varma, J. Nayak, K. G. Suresh, and A. Alam, FeRhCrSi: Spin semimetal with spin-valve behavior at room temperature, Phys. Rev. B 107, L100401 (2023).
- Nishihaya et al. [2025] S. Nishihaya, M. J. A. Jardine, H. S. Inbar, A. Goswami, J. T. Dong, A. N. Engel, Y.-H. Chang, C. P. Dempsey, M. Hashimoto, D. Lu, N. Marom, and C. J. Palmstrøm, Near-half-metallic state in the half-Heusler PtMnSb film on a III-V substrate, Phys. Rev. Mater. 9, 044403 (2025).
- de Groot et al. [1983] R. A. de Groot, F. M. Mueller, P. G. v. Engen, and K. H. J. Buschow, New class of materials: Half-metallic ferromagnets, Phys. Rev. Lett. 50, 2024 (1983).
- Žutić et al. [2004] I. Žutić, J. Fabian, and S. Das Sarma, Spintronics: Fundamentals and applications, Rev. Mod. Phys. 76, 323 (2004).
- Katsnelson et al. [2008] M. I. Katsnelson, V. Y. Irkhin, L. Chioncel, A. I. Lichtenstein, and R. A. de Groot, Half-metallic ferromagnets: From band structure to many-body effects, Rev. Mod. Phys. 80, 315 (2008).
- Graf et al. [2011] T. Graf, C. Felser, and S. S. Parkin, Simple rules for the understanding of Heusler compounds, Progress in Solid State Chemistry 39, 1 (2011).
- Husmann and Singh [2006] A. Husmann and L. J. Singh, Temperature dependence of the anomalous hall conductivity in the Heusler alloy , Phys. Rev. B 73, 172417 (2006).
- Shigeta et al. [2018] I. Shigeta, Y. Fujimoto, R. Ooka, Y. Nishisako, M. Tsujikawa, R. Y. Umetsu, A. Nomura, K. Yubuta, Y. Miura, T. Kanomata, M. Shirai, J. Gouchi, Y. Uwatoko, and M. Hiroi, Pressure effect on the magnetic properties of the half-metallic Heusler alloy , Phys. Rev. B 97, 104414 (2018).
- Kübler et al. [2007] J. Kübler, G. H. Fecher, and C. Felser, Understanding the trend in the curie temperatures of -based Heusler compounds: Ab initio calculations, Phys. Rev. B 76, 024414 (2007).
- Fortunato et al. [2024] N. M. Fortunato, X. Li, S. Schönecker, R. Xie, A. Taubel, F. Scheibel, I. Opahle, O. Gutfleisch, and H. Zhang, High-throughput screening of all--metal Heusler alloys for magnetocaloric applications, Chemistry of Materials 36, 6765 (2024).
- Marathe and Herper [2023] M. Marathe and H. C. Herper, Exploration of all- Heusler alloys for permanent magnets: An ab initio based high-throughput study, Phys. Rev. B 107, 174402 (2023).
- Guo et al. [2025] Y. Guo, C. Guo, Y. Xu, B. Zheng, X. Ni, and F. Tian, Machine learning investigation of chemical disorder effects in 24-valence-electron full-Heusler alloys with single magnetic element, Journal of Magnetism and Magnetic Materials 629, 173322 (2025).
- Picozzi et al. [2004] S. Picozzi, A. Continenza, and A. J. Freeman, Role of structural defects on the half-metallic character of and Heusler alloys, Phys. Rev. B 69, 094423 (2004).
- Devi et al. [2018] P. Devi, S. Singh, B. Dutta, K. Manna, S. W. D’Souza, Y. Ikeda, E. Suard, V. Petricek, P. Simon, P. Werner, S. Chadhov, S. S. P. Parkin, C. Felser, and D. Pandey, Adaptive modulation in the magnetic shape-memory Heusler alloy, Phys. Rev. B 97, 224102 (2018).
- Kanomata et al. [2010] T. Kanomata, Y. Chieda, K. Endo, H. Okada, M. Nagasako, K. Kobayashi, R. Kainuma, R. Y. Umetsu, H. Takahashi, Y. Furutani, H. Nishihara, K. Abe, Y. Miura, and M. Shirai, Magnetic properties of the half-metallic Heusler alloys and under pressure, Phys. Rev. B 82, 144415 (2010).
- Suto et al. [2024] H. Suto, V. Barwal, K. Simalaotao, Z. Li, K. Masuda, T. Sasaki, Y. Miura, and Y. Sakuraba, Negative spin polarization of Mn2VGa Heusler alloy thin films studied in current-perpendicular-to-plane giant magnetoresistance devices, Journal of Applied Physics 135, 203901 (2024).
- Ouardi et al. [2013] S. Ouardi, G. H. Fecher, C. Felser, and J. Kübler, Realization of spin gapless semiconductors: The Heusler compound , Phys. Rev. Lett. 110, 100401 (2013).
- Životský et al. [2022] O. Životský, K. Skotnicová, T. Čegan, J. Juřica, L. Gembalová, F. Zažímal, and I. Szurman, Structural and magnetic properties of inverse-Heusler Mn2FeSi alloy powder prepared by ball milling, Materials 15, 697 (2022).
- Wurmehl et al. [2005] S. Wurmehl, G. H. Fecher, H. C. Kandpal, V. Ksenofontov, C. Felser, H.-J. Lin, and J. Morais, Geometric, electronic, and magnetic structure of : Curie temperature and magnetic moment measurements and calculations, Phys. Rev. B 72, 184434 (2005).
- Singh and Mazin [1998] D. J. Singh and I. I. Mazin, Electronic structure, local moments, and transport in , Phys. Rev. B 57, 14352 (1998).
- Khmelevskyi et al. [2015] S. Khmelevskyi, E. Simon, and L. Szunyogh, Antiferromagnetism in full Heusler alloys: Effects of magnetic frustration and chemical disorder, Phys. Rev. B 91, 094432 (2015).
- Sokolovskiy et al. [2012] V. V. Sokolovskiy, V. D. Buchelnikov, M. A. Zagrebin, P. Entel, S. Sahoo, and M. Ogura, First-principles investigation of chemical and structural disorder in magnetic Ni2Mn1+xSn1-x Heusler alloys, Phys. Rev. B 86, 134418 (2012).
- Decolvenaere et al. [2019] E. Decolvenaere, E. Levin, R. Seshadri, and A. Van der Ven, Modeling magnetic evolution and exchange hardening in disordered magnets: The example of Heusler alloys, Phys. Rev. Mater. 3, 104411 (2019).
- Li et al. [2013] T. Li, J. Duan, C. Yang, and X. Kou, Synthesis, microstructure and magnetic properties of Heusler Co2FeSn nanoparticles, Micro & Nano Letters 8, 143 (2013).
- Smith et al. [2024] R. Smith, Z. Gercsi, R. Zhang, K. Siewierska, K. Rode, and J. Coey, Effects of disorder on the magnetic properties of the Heusler alloy V2FeAl, Acta Materialia 267, 119733 (2024).
- Miroshkina et al. [2022] O. N. Miroshkina, B. Eggert, J. Lill, B. Beckmann, D. Koch, M. Y. Hu, T. Lojewski, S. Rauls, F. Scheibel, A. Taubel, M. Šob, K. Ollefs, O. Gutfleisch, H. Wende, M. E. Gruner, and M. Friák, Impact of magnetic and antisite disorder on the vibrational densities of states in Heusler alloys, Phys. Rev. B 106, 214302 (2022).
- Galanakis et al. [2002] I. Galanakis, P. H. Dederichs, and N. Papanikolaou, Slater-Pauling behavior and origin of the half-metallicity of the full-Heusler alloys, Phys. Rev. B 66, 174429 (2002).
- Chico et al. [2016] J. Chico, S. Keshavarz, Y. Kvashnin, M. Pereiro, I. Di Marco, C. Etz, O. Eriksson, A. Bergman, and L. Bergqvist, First-principles studies of the Gilbert damping and exchange interactions for half-metallic Heuslers alloys, Phys. Rev. B 93, 214439 (2016).
- Kuroda et al. [2020] F. Kuroda, T. Fukushima, and T. Oguchi, First-principles study of magnetism and phase stabilities of V2 based antiferromagnetic Heusler alloys, Journal of Applied Physics 127, 193904 (2020).
- Skaftouros et al. [2013] S. Skaftouros, K. Özdoğan, E. Şaşıoğlu, and I. Galanakis, Generalized Slater-Pauling rule for the inverse Heusler compounds, Phys. Rev. B 87, 024420 (2013).
- Kresse and Furthmüller [1996] G. Kresse and J. Furthmüller, Efficient iterative schemes for ab initio total-energy calculations using a plane-wave basis set, Phys. Rev. B 54, 11169 (1996).
- Kresse and Joubert [1999] G. Kresse and D. Joubert, From ultrasoft pseudopotentials to the projector augmented-wave method, Phys. Rev. B 59, 1758 (1999).
@@
- Blöchl [1994] P. E. Blöchl, Projector augmented-wave method, Phys. Rev. B 50, 17953 (1994).
- Perdew et al. [1996] J. P. Perdew, K. Burke, and M. Ernzerhof, Generalized gradient approximation made simple, Phys. Rev. Lett. 77, 3865 (1996).
- Hohenberg and Kohn [1964] P. Hohenberg and W. Kohn, Inhomogeneous electron gas, Phys. Rev. 136, B864 (1964).
- Kohn and Sham [1965] W. Kohn and L. J. Sham, Self-consistent equations including exchange and correlation effects, Phys. Rev. 140, A1133 (1965).
- Li et al. [2016] P. Li, X. Liu, M. Chen, P. Lin, X. Ren, L. Lin, C. Yang, and L. He, Large-scale ab initio simulations based on systematically improvable atomic basis, Computational Materials Science 112, 503 (2016), computational Materials Science in China.
- Zhou et al. [2025] W. Zhou, D. Zheng, Q. Liu, D. Lu, Y. Liu, P. Lin, Y. Huang, X. Peng, J. J. Bao, C. Cai, Z. Jin, J. Wu, H. Zhang, G. Jin, Y. Ji, Z. Shen, X. Liu, L. Sun, Y. Cao, M. Sun, J. Liu, T. Chen, R. Liu, Y. Li, H. Han, X. Liang, T. Bao, N. Chen, H. Ren, X. Zhang, Z. Liu, Y. Fu, M. Liu, Z. Li, T. Wen, Z. Tang, Y. Xu, W. Duan, X. Wang, Q. Gu, F.-Z. Dai, Q. Zheng, J. Zhao, Y. Zhang, Q. Ou, H. Jiang, S. Liu, B. Xu, S. Xu, X. Ren, L. He, L. Zhang, and M. Chen, ABACUS: An electronic structure analysis package for the AI era, arXiv e-prints , 2501.08697 (2025).
- Schlipf and Gygi [2015] M. Schlipf and F. Gygi, Optimization algorithm for the generation of ONCV pseudopotentials, Computer Physics Communications 196, 36 (2015).
- He et al. [2021] X. He, N. Helbig, M. J. Verstraete, and E. Bousquet, TB2J: A python package for computing magnetic interaction parameters, Computer Physics Communications 264, 107938 (2021).
- Skubic et al. [2008] B. Skubic, J. Hellsvik, L. Nordström, and O. Eriksson, A method for atomistic spin dynamics simulations: implementation and examples, Journal of Physics: Condensed Matter 20, 315203 (2008).
- Goraus et al. [2023] J. Goraus, M. Fijałkowski, J. Czerniewski, and W. Gumulak, Magnetic properties of V2MnGa, V2MnAl, V2FeGa and V2FeAl, Journal of Magnetism and Magnetic Materials 585, 171103 (2023).
- Zhang et al. [2012] X. Zhang, X. Dai, G. Chen, H. Liu, H. Luo, Y. Li, W. Wang, G. Wu, and G. Liu, First-principle study for full-Heusler compounds V2YAl (Y=V, Cr, Mn, Fe, Co, Ni) and a discussion to Slater–Pauling rule, Computational Materials Science 59, 1 (2012).
@@
- Faleev et al. [2017] S. V. Faleev, Y. Ferrante, J. Jeong, M. G. Samant, B. Jones, and S. S. P. Parkin, Unified explanation of chemical ordering, the Slater-Pauling rule, and half-metallicity in full Heusler compounds, Phys. Rev. B 95, 045140 (2017).
- Xu et al. [2024] Y. Xu, Y.-C. Wang, X. Jin, H. Liu, Y. Liu, H. Song, and F. Tian, Mechanism of magnetic phase transition in correlated magnetic metal: insight into itinerant ferromagnet Fe3‑δGeTe2, Commun. Phys. 7, 381 (2024).
@@
- Xu et al. [2025a] Y. Xu, X. Jin, J. Xiang, H. Zhang, and F. Tian, investigation of the magnetic phase transition in the itinerant ferromagnet , Phys. Rev. B 111, 155142 (2025a).
- Xu et al. [2025b] Y. Xu, X. Jin, H. Tang, and F. Tian, Electronic structures and magnetism in van der waals flat-band material , Europhysics Letters 151, 36002 (2025b).

### v2

@@ III.4 Magnetic transition temperature of - and -V2Al
The temperature dependence of magnetic phase transitions in -type VAl compounds differs markedly from that in -type structures. As shown in Fig.4(b), the Curie temperatures in -type compounds vary within a narrower range, from approximately 700 K to 300 K, compared to the broader variation in -type materials. This narrower range reflects the more complex network of magnetic exchange couplings in the inverse-Heusler structure. Specifically, V2CrAl exhibits the highest at about 700 K, while the Curie temperature gradually decreases as changes from Fe to Ni. The magnetic transition temperature of V2MnAl shows a slight anomaly, lying between the values for Co and Ni. Despite the apparent complexity of these trends, they can be rationalized using the proposed microscopic mechanism based on the triangular exchange pathways. As discussed above, the NN exchange coupling involving the -site dominates the magnetic transition behavior. For instance, the relatively large contributes to the higher in V2CrAl. Although is also sizable, the of V2MnAl is significantly lower due to the additional influence of the coupling. In V2CrAl, the ferromagnetic – interactions enhance , whereas in V2MnAl, the coupling is antiferromagnetic, leading to a notable reduction of . In the series Fe, Co, Ni, the gradual decrease of is primarily driven by the weakening the absolute value of the NN coupling, which diminishes from to 0.1 meV (see Supplementary Information).
@@
III.5 Effect of chemical disorder on magnetism of partially-ordered V2Al
@@
Our calculations on partially chemically disordered VAl (for Cr, Co, Ni, see Supplementary Information) indicate that the NN magnetic exchange coupling remains the dominant factor governing the magnetic transition temperature. Notably, during the transition from the to the -type structure, the 4 site is consistently occupied by V2 atoms, ensuring that the V-X-V magnetic triangular pathway is preserved. Consequently, the microscopic magnetic mechanism in chemically disordered VAl is still primarily governed by the V-X-V triangular pathway and the associated NN magnetic exchange couplings .
III.6 Electronic structures of fully-ordered V2Al
Based on the calculated magnetic ground states of -type VAl in Sec. III.2, the observed complex magnetic orders and non-monotonic variation of total magnetic moments with atomic composition are proposed related to their electronic structures. In -type VAl, the relatively small total magnetic moments of V2CrAl and V2MnAl originate from the anti-parallel alignment of spin moments between the V and sublattices. This behavior is clearly reflected in the calculated density of states (DOS). As shown in Fig. 7(a), the total DOS of -type V2Al exhibits a high value at the Fermi level, indicating itinerant magnetism in these compounds [54, 53, 52]. Notably, several compounds display nearly half-metallic character. As shown in Fig. 7(b), Cr and Mn atoms exhibit pronounced occupied states with sharp peaks at approximately -1 eV and -2 eV in the spin-down channel, while their spin-up channel states are minimally occupied. In contrast, for Fe, Co, and Ni, the PDOS of atoms shows significant occupation in the spin-up channel and comparatively low values in the spin-down channel. These electronic features are consistent with the observed transition in magnetic order from ferrimagnetism ( Cr, Mn) to ferromagnetism ( Fe, Co, Ni).
@@ III.6 Electronic structures of fully-ordered V2Al
From our calculated electronic structures, we observe substantial hybridization among the V-, -, and Al- states near the Fermi level in all compounds, as evidenced by the similar peak shapes and their alignment in energy. Although such hybridization is an evident feature throughout the VAl series, the key factor governing the evolution of magnetic behavior is the variation in band filling associated with the -site element. A comparison of the DOS for Cr and Mn with those for Fe, Co, and Ni clearly shows that, in the Cr and Mn systems, the -electron manifold lies much closer to half-filling. This near half-filled configuration enhances antiferromagnetic superexchange between V and Cr (Mn) atoms through virtual electron-hopping processes between neighboring sites [41, 38]. In contrast, as the -site element becomes more electron-rich (from Fe to Ni), the increased band filling shifts the DOS away from half-filling, thereby suppressing the antiferromagnetic superexchange channel and promoting ferromagnetic exchange.
@@ IV CONCLUSION
In this work, we employed density functional theory calculations combined with Monte Carlo simulations based on Heisenberg magnetic exchange couplings to investigate V2Al alloys and introduced the concept of “magnetic motifs”. These magnetic motifs govern both the magnetic ground states and the magnetic transition temperatures in fully- and partially-ordered compounds. The underlying physical mechanism arises from the modulation of NN magnetic exchange couplings within the V--V triangular pathways, which is influenced by atomic composition, crystal structure, and the degree of chemical disorder. Using the framework of magnetic motifs, the magnetic orders can be rationalized in terms of the V--V triangular pathways formed by NN couplings. Furthermore, the magnetic transition temperatures are predominantly determined by both NN and additional magnetic exchange interactions, irrespective of whether the alloy is fully or partially ordered. The concept of “magnetic motifs” captures the minimal set of dominant exchange interactions that governs both the magnetic ground states and the evolution of the Curie temperature across different structures and compositions. We anticipate that identifying such magnetic motifs in other Heusler alloys will provide a systematic and predictive framework for understanding and tailoring their diverse magnetic properties. Furthermore, a comprehensive investigation that combines phase stability analysis with detailed magnetic characterization will be essential for guiding the design of high-performance Heusler-based spintronic devices in the future studies.
@@ Acknowledgements.
This work is supported by the National Natural Science Foundation of China (Grants No. 52371174 and No. 12204033), Science Challenge Project (Grant No.TZ2025009), the State Key Lab of Advanced Metals and Materials (Grant No. 2025Z-Z22). Numerical computations were performed at the High-performance Computing Platform of University of Science and Technology Beijing.References
- [1] Note: See Supplemental Material at [URL will be inserted by publisher] for more theoretical analysis. Cited by: §II.
@@
- [2] (1994-12) Projector augmented-wave method. Phys. Rev. B 50, pp. 17953–17979. External Links: Document, Link Cited by: §II.
- [3] (2016-06) First-principles studies of the Gilbert damping and exchange interactions for half-metallic Heuslers alloys. Phys. Rev. B 93, pp. 214439. External Links: Document, Link Cited by: §I.
- [4] (1983-06) New class of materials: half-metallic ferromagnets. Phys. Rev. Lett. 50, pp. 2024–2027. External Links: Document, Link Cited by: §I.
- [5] (2019-10) Modeling magnetic evolution and exchange hardening in disordered magnets: the example of Heusler alloys. Phys. Rev. Mater. 3, pp. 104411. External Links: Document, Link Cited by: §I.
- [6] (2018-06) Adaptive modulation in the magnetic shape-memory Heusler alloy. Phys. Rev. B 97, pp. 224102. External Links: Document, Link Cited by: §I.
@@
- [7] (2017-01) Unified explanation of chemical ordering, the Slater-Pauling rule, and half-metallicity in full Heusler compounds. Phys. Rev. B 95, pp. 045140. External Links: Document, Link Cited by: §III.2.
- [8] (2024) High-throughput screening of all--metal Heusler alloys for magnetocaloric applications. Chem. Mater. 36 (14), pp. 6765–6776. External Links: Document, Link Cited by: §I.
- [9] (2002-11) Slater-Pauling behavior and origin of the half-metallicity of the full-Heusler alloys. Phys. Rev. B 66, pp. 174429. External Links: Document, Link Cited by: §I, §III.1.
- [10] (2023) Magnetic properties of V2MnGa, V2MnAl, V2FeGa and V2FeAl. J. Magn. Magn. Mater. 585, pp. 171103. External Links: ISSN 0304-8853, Document, Link Cited by: Table 1, Table 1, Table 1, §III.2, §III.2, Table 2, Table 2.
- [11] (2011) Simple rules for the understanding of Heusler compounds. Prog. Solid State Chem. 39 (1), pp. 1–50. External Links: ISSN 0079-6786, Document, Link Cited by: §I, §III.1.
- [12] (2025) Machine learning investigation of chemical disorder effects in 24-valence-electron full-Heusler alloys with single magnetic element. J. Magn. Magn. Mater. 629, pp. 173322. External Links: ISSN 0304-8853, Document, Link Cited by: §I.
- [13] (2023-07) High spin polarization in the disordered quaternary heusler alloy femnvga. Phys. Rev. B 108, pp. 045137. External Links: Document, Link Cited by: §I.
- [14] (2021) TB2J: a python package for computing magnetic interaction parameters. Comput. Phys. Commun. 264, pp. 107938. External Links: ISSN 0010-4655, Document, Link Cited by: §II.
- [15] (1964-11) Inhomogeneous electron gas. Phys. Rev. 136, pp. B864–B871. External Links: Document, Link Cited by: §II.
- [16] (2006-05) Temperature dependence of the anomalous hall conductivity in the Heusler alloy . Phys. Rev. B 73, pp. 172417. External Links: Document, Link Cited by: §I.
- [17] (2010-10) Magnetic properties of the half-metallic Heusler alloys and under pressure. Phys. Rev. B 82, pp. 144415. External Links: Document, Link Cited by: §I.
- [18] (2008-04) Half-metallic ferromagnets: from band structure to many-body effects. Rev. Mod. Phys. 80, pp. 315–378. External Links: Document, Link Cited by: §I.
- [19] (2015-03) Antiferromagnetism in full Heusler alloys: effects of magnetic frustration and chemical disorder. Phys. Rev. B 91, pp. 094432. External Links: Document, Link Cited by: §I.
- [20] (2015-12) The Open Quantum Materials Database (OQMD): assessing the accuracy of DFT formation energies. npj Computational Materials 1 (1), pp. 15010. External Links: Document, Link Cited by: §III.1.
- [21] (1965-11) Self-consistent equations including exchange and correlation effects. Phys. Rev. 140, pp. A1133–A1138. External Links: Document, Link Cited by: §II.
- [22] (1996-10) Efficient iterative schemes for ab initio total-energy calculations using a plane-wave basis set. Phys. Rev. B 54, pp. 11169–11186. External Links: Document, Link Cited by: §II.
- [23] (1999-01) From ultrasoft pseudopotentials to the projector augmented-wave method. Phys. Rev. B 59, pp. 1758–1775. External Links: Document, Link Cited by: §II.
- [24] (2007-07) Understanding the trend in the Curie temperatures of -based Heusler compounds: Ab initio calculations. Phys. Rev. B 76, pp. 024414. External Links: Document, Link Cited by: §II.
- [25] (2007-07) Understanding the trend in the curie temperatures of -based Heusler compounds: ab initio calculations. Phys. Rev. B 76, pp. 024414. External Links: Document, Link Cited by: §I.
- [26] (2020-05) First-principles study of magnetism and phase stabilities of V2 based antiferromagnetic Heusler alloys. J. Appl. Phys. 127 (19), pp. 193904. External Links: ISSN 0021-8979, Document, Link Cited by: §I.
- [27] (2016) Large-scale ab initio simulations based on systematically improvable atomic basis. Comput. Mater. Sci. 112, pp. 503–517. Note: Computational Materials Science in China External Links: ISSN 0927-0256, Document, Link Cited by: §II.
- [28] (2013) Synthesis, microstructure and magnetic properties of Heusler Co2FeSn nanoparticles. Micro & Nano Lett. 8 (3), pp. 143–146. External Links: Document, Link Cited by: §I.
- [29] (2023-10) Doubly Screened Coulomb Correction approach for strongly correlated systems. J. Phys. Chem. Lett. 14, pp. 8930–8939. External Links: Document, Link Cited by: §II.
- [30] (2023-05) Exploration of all- Heusler alloys for permanent magnets: an ab initio based high-throughput study. Phys. Rev. B 107, pp. 174402. External Links: Document, Link Cited by: §I.
- [31] (2022-12) Impact of magnetic and antisite disorder on the vibrational densities of states in Heusler alloys. Phys. Rev. B 106, pp. 214302. External Links: Document, Link Cited by: §I.
- [32] (2024) Orbital-engineered anomalous Hall conductivity in stable full Heusler compounds: a pathway to optimized spintronics. Comput. Mater. Sci. 12 (37), pp. 15082–15091. External Links: Document Cited by: §III.1.
- [33] (2025-04) Near-half-metallic state in the half-Heusler PtMnSb film on a III-V substrate. Phys. Rev. Mater. 9, pp. 044403. External Links: Document, Link Cited by: §I.
- [34] (2013-03) Realization of spin gapless semiconductors: the Heusler compound . Phys. Rev. Lett. 110, pp. 100401. External Links: Document, Link Cited by: §I.
- [35] (2001-10) Ab initio calculations of exchange interactions, spin-wave stiffness constants, and Curie temperatures of Fe, Co, and Ni. Phys. Rev. B 64, pp. 174402. External Links: Document, Link Cited by: §II.
- [36] (1996-10) Generalized gradient approximation made simple. Phys. Rev. Lett. 77, pp. 3865–3868. External Links: Document, Link Cited by: §II.
- [37] (2004-03) Role of structural defects on the half-metallic character of and Heusler alloys. Phys. Rev. B 69, pp. 094423. External Links: Document, Link Cited by: §I, §III.1.
- [38] (2013-04) Antiferromagnetic exchange interactions in the Ni2Mn1.4In0.6 ferromagnetic Heusler alloy. Phys. Rev. B 87, pp. 144412. External Links: Document, Link Cited by: §III.6.
- [39] (2013-11) Materials design and discovery with high-throughput density functional theory: the Open Quantum Materials Database (OQMD). JOM 65 (11), pp. 1501–1509. External Links: Document, Link Cited by: §III.1.
- [40] (2005-11) Exchange interactions and temperature dependence of magnetization in half-metallic Heusler alloys. Phys. Rev. B 72, pp. 184415. External Links: Document, Link Cited by: §II.
- [41] (2008-02) Role of conduction electrons in mediating exchange interactions in Mn-based Heusler alloys. Phys. Rev. B 77, pp. 064417. External Links: Document, Link Cited by: §III.6.
- [42] (2015) Optimization algorithm for the generation of ONCV pseudopotentials. Comput. Phys. Commun. 196, pp. 36–44. External Links: ISSN 0010-4655, Document, Link Cited by: §II.
- [43] (2018-03) Pressure effect on the magnetic properties of the half-metallic Heusler alloy . Phys. Rev. B 97, pp. 104414. External Links: Document, Link Cited by: §I.
- [44] (1998-06) Electronic structure, local moments, and transport in . Phys. Rev. B 57, pp. 14352–14356. External Links: Document, Link Cited by: §I.
- [45] (2013-01) Generalized Slater-Pauling rule for the inverse Heusler compounds. Phys. Rev. B 87, pp. 024420. External Links: Document, Link Cited by: §II, §III.2, Table 2.
- [46] (2008-07) A method for atomistic spin dynamics simulations: implementation and examples. J. Phys.: Condens. Matter 20 (31), pp. 315203. External Links: Document, Link Cited by: §II.
- [47] (2024) Effects of disorder on the magnetic properties of the Heusler alloy V2FeAl. Acta Mater. 267, pp. 119733. External Links: ISSN 1359-6454, Document, Link Cited by: §I, §III.1.
- [48] (2012-10) First-principles investigation of chemical and structural disorder in magnetic Ni2Mn1+xSn1-x Heusler alloys. Phys. Rev. B 86, pp. 134418. External Links: Document, Link Cited by: §I.
- [49] (2024-05) Negative spin polarization of Mn2VGa Heusler alloy thin films studied in current-perpendicular-to-plane giant magnetoresistance devices. J. Appl. Phys. 135 (20), pp. 203901. External Links: ISSN 0021-8979, Document, Link Cited by: §I.
- [50] (2023-03) FeRhCrSi: spin semimetal with spin-valve behavior at room temperature. Phys. Rev. B 107, pp. L100401. External Links: Document, Link Cited by: §I.
- [51] (2005-11) Geometric, electronic, and magnetic structure of : curie temperature and magnetic moment measurements and calculations. Phys. Rev. B 72, pp. 184434. External Links: Document, Link Cited by: §I.
- [52] (2025-08) Electronic structures and magnetism in van der waals flat-band material . Europhys. Lett. 151 (3), pp. 36002. External Links: Document, Link Cited by: §III.6.
@@
- [53] (2025-04) Investigation of the magnetic phase transition in the itinerant ferromagnet . Phys. Rev. B 111, pp. 155142. External Links: Document, Link Cited by: §III.6.
- [54] (2024-11) Mechanism of magnetic phase transition in correlated magnetic metal: insight into itinerant ferromagnet Fe3‑δGeTe2. Commun. Phys. 7 (1), pp. 381. External Links: Document Cited by: §III.6.
- [55] (2012) First-principle study for full-Heusler compounds V2YAl (Y=V, Cr, Mn, Fe, Co, Ni) and a discussion to Slater–Pauling rule. Comput. Mater. Sci. 59, pp. 1–5. External Links: ISSN 0927-0256, Document, Link Cited by: §II, §III.1, §III.2, Table 2, Table 2.
- [56] (2025-01) ABACUS: an electronic structure analysis package for the AI era. arXiv, pp. 2501.08697. External Links: Document Cited by: §II.
- [57] (2022) Structural and magnetic properties of inverse-Heusler Mn2FeSi alloy powder prepared by ball milling. Materials 15 (3), pp. 697. External Links: Link, ISSN 1996-1944, Document Cited by: §I.
- [58] (2004-04) Spintronics: fundamentals and applications. Rev. Mod. Phys. 76, pp. 323–410. External Links: Document, Link Cited by: §I.